# Physics of droplet regulation in biological cells

### v3

@@ Abstract
  • Version of draft: September 12, 2025
@@ 2.1.1 Thermodynamics of liquids
where the first term accounts for the mixing entropy with $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$ due to incompressibility; see appendix A. We also use the molecular volume $\nu_{0}$ of the solvent to define non-dimensional molecular sizes $\Omega_{i}=\nu_{i}/\nu_{0}$. The last two terms in equation (2.1) specify enthalpic interactions using non-dimensional quantities $w_{i}$ and $\chi_{ij}$, which include solvent interactions (see equation (A.11)) and whose particular values can be determined from detailed theories [85]. Here, $w_{i}$ captures the difference of internal energy between species $i$ and the solvent, but it can also account for external potentials. Such potentials can in principle depend on space, due to, for example, gravity or gradients in a regulatory species [24], but we do not discuss this aspect further in this review. In contrast, the symmetric Flory parameters $\chi_{ij}=\chi_{ji}$ capture interactions: positive values raise the free energy and thus relate to repulsion between species $i$ and $j$, while negative values imply effective attraction. Since the interactions $\chi_{ij}$ are measured relative to solvent molecules, the diagonal terms need not vanish; see appendix A. The Flory–Huggins free energy given by equation (2.1) only describes the basic behavior of phase separating mixtures, and we discuss more complex interactions in section 3.1.
@@ 2.1.2 Kinetics of heterogeneous systems
We now derive kinetic equations based on linear non-equilibrium thermodynamics [110, 101], starting from the continuity equation, $\partial_{t}\phi_{i}+\nabla\cdot\boldsymbol{J}_{i}=s_{i}$, where $\boldsymbol{J}_{i}$ denotes spatial fluxes and the source term $s_{i}(\boldsymbol{r})$ accounts for chemical reactions as described in section 5. To separate fluxes due to hydrodynamic flows (associated with momentum transport) from diffusive fluxes driven by gradients in chemical potentials, we split $\boldsymbol{J}_{i}$ into a centre-of-mass velocity field $\boldsymbol{v}=\sum_{i=0}^{N_{\mathrm{c}}}\boldsymbol{J}_{i}$ and relative fluxes $\boldsymbol{j}_{i}=\phi_{i}(\boldsymbol{v}_{i}-\boldsymbol{v})$, where $\boldsymbol{v}_{i}=\boldsymbol{J}_{i}/\phi_{i}$ is the velocity of component $i$ in the lab frame. Since many biomolecules have mass densities close to water, we here assumed equal mass density for all components, and refer the reader to ref. [101] for the general case. Hence,
@@
$\mathcal{J}_{ij}$ $=-\nu_{0}\Omega_{i}^{2}\boldsymbol{q}^{2}\sum_{l=1}^{N_{\mathrm{c}}}\Lambda_{il}\left[\frac{\partial^{2}f}{\partial\phi_{l}\phi_{j}}+\kappa_{lj}\boldsymbol{q}^{2}\right]+\frac{\partial s_{i}}{\partial\phi_{j}}$ (2.15)
has only negative eigenvalues for all wave vectors $\boldsymbol{q}$. In contrast, the homogeneous state is unstable if $\mathcal{J}_{ij}$ has positive eigenvalues and the largest eigenvalue will typically determine the dominant length scale $2\pi/|\boldsymbol{q}|$ from the associated wavevector $\boldsymbol{q}$. However, since equations (2.14) are nonlinear, such a stability analysis only provides information about the initial dynamics.
@@ 2.1.3 Behaviour of binary liquids
Figure 2.1: Phase behaviour of a binary liquid. (A) Free energy density $f$ given by equation (2.16) as a function of the fraction $\phi$ (blue line) for $\Omega=1$, $\chi=2.5$, and $w=0.2$. The common-tangent construction (black line) solves equations (2.30) and thus determines the equilibrium fractions $\phi_{\mathrm{out}}$ and $\phi_{\mathrm{in}}$ (black disks) without Laplace pressure $P_{\gamma}$. Finite $P_{\gamma}$ require two parallel tangents offset by $P_{\gamma}$ (gray lines). The inflection points (orange disks) enclose the spinodal region where $f^{\prime\prime}(\phi)<0$. (B) Exchange chemical potential $\bar{\mu}$ given by equation (2.17) as a function of $\phi$ corresponding to panel A. The blue regions of equal area indicating a Maxwell construction, equivalent to the common-tangent construction. (C) Osmotic pressure $\Pi=\nu^{-1}\phi\bar{\mu}-f$ as a function of $\phi$ corresponding to panel A. (D) $\Pi$ as a function of $\bar{\mu}$ corresponding to panel A. (E, F) Phase diagrams corresponding to panel A highlighting the equilibrium volume fractions (black binodal line) and the spinodal region (within the orange spinodal line) for a symmetric mixture ($\Omega=1$, panel E) and for large solutes ($\Omega=10$, panel F). The critical point is marked by a star and grey areas denotes the region where bi-continuous structures have lower energy than droplets in two dimensions (obtained by comparing interfacial energies of disks and stripes for equal length scales). (G, H) Grand-canonical phase diagram corresponding to panels E and F. The binodals (black lines) mark the first-order transition between dilute and dense phases. Both phases are linearly stable in the spinodal region (between orange dotted lines).
In many cases, we do not need to describe all the individual components of a mixture and can instead focus on droplet material separating from a solvent. In this case, the mixture is treated as a binary fluid, and only the volume fraction $\phi$ of one component needs to be specified (${N_{\mathrm{c}}}=1$}). The local free energy density then reduces to
@@
The resulting spinodal line $\chi_{\mathrm{spin}}(\phi)$ encloses a region in phase space, see figure 2.1. Within this region the homogeneous state is unstable and develops patterns at a dominant length scale $l_{\mathrm{max}}=\pi[-8\kappa/f^{\prime\prime}(\phi)]^{1/2}$ that eventually lead to phase separation, which is known as spinodal decomposition [24]. The minimum of the curve $\chi_{\mathrm{spin}}(\phi)$ corresponds to the critical point,
@@ 2.1.4 Interfaces and phase morphologies
In the absence of chemical reactions ($s=0$), the dynamics described by equation (2.23) result in equilibrium states that minimise the free energy $F$ given by equation (2.22). Figure 2.2A shows that the corresponding equilibrium profile $\phi_{\mathrm{eq}}(x)$ smoothly interpolates between the fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ of the two co-existing phases. Denoting by $x$ a coordinate perpendicular to the interface positioned at $x=0$, the profile can be approximted as
$\phi_{\mathrm{eq}}(x)$ $\approx\frac{\phi^{(0)}_{\mathrm{in}}+\phi^{(0)}_{\mathrm{out}}}{2}+\frac{\phi^{(0)}_{\mathrm{in}}-\phi^{(0)}_{\mathrm{out}}}{2}\tanh\left(\frac{2x}{0pt}\right)\;,$ (2.24)
where $0pt$ denotes the interfacial width. This expression is exact for a simpler polynomial free energy [134, 24]. For our case and symmetric molecular sizes ($\Omega=1$), the interfacial width $0pt$ is approximately (appendix C)
@@
and the sigmoidal profile given by equation (2.24) provides a good approximation.
In the typical case where the region of the interface is thin compared to the linear size of the adjacent phases, one can use a thin-interface approximation and replace the sigmoidal profile by a step function. The excess free energy, which is not accounted for by the bulk phases, defines the surface energy
@@
where $\bar{\mu}_{\mathrm{eq}}$ is the exchange chemical potential and $\Pi_{\mathrm{eq}}=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{out}}-f(\phi^{(0)}_{\mathrm{out}})=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{in}}-f(\phi^{(0)}_{\mathrm{in}})$ denotes the associated osmotic pressure; see appendix C. The second line assumes a monotonic profile $\phi_{\mathrm{eq}}(x)$, which may not be valid in complex mixtures [135]. However, $\phi_{\mathrm{eq}}(x)$ is always monotonic in binary mixtures, and in the symmetric case $\Omega=1$, we find that
$\gamma$ $\approx\sqrt{\frac{k_{\mathrm{B}}T\kappa}{2\nu}}\cdot\frac{\phi^{(0)}_{\mathrm{out}}-\phi^{(0)}_{\mathrm{in}}}{3}\ln\left(4\phi^{(0)}_{\mathrm{in}}\phi^{(0)}_{\mathrm{out}}\right)\;$; (2.27)
@@
The thin-interface approximation simplifies the analysis of equilibrium states in binary liquids: material conservation fixes the total volume of the dilute and dense phases, so their shape is the only degree of freedom. The total free energy thus comprises a constant term covering the bulk contributions and an interfacial term, $\gamma A$, where $A$ is the total interfacial area. Equilibrium states thus minimise $A$, which implies constant mean curvature of interfaces [138]. Consequently, surface energies can be interpreted as a mechanical surface tension [139], explaining why the minority phase favours spherical shapes. Conversely, if both phases occupy similar volumes, bicontinuous structures, such as cylinders and flat lamella, can emerge; see figure 2.2. These structures predominate in a region of phase space that overlaps significantly with the spinodal region; see figure 2.1E,F. However, these two phenomena are distinct, so droplets can for instance emerge from spinodal decomposition. Bicontinuous structures are rarely observed in biological cells, and so we exclusively focus on droplets in this review.
@@ 2.2.1 Local equilibrium at the interface
where $c_{\mathrm{in}}=\phi^{(0)}_{\mathrm{in}}/\nu\Omega$ is the number concentration inside the droplet whereas the dilute phase is approximated as an ideal solution. When only surface tension acts on the droplet, $\Delta P=P_{\gamma}$, we find the Gibbs-Thomson relation [142]
@@
assuming the radius $R$ is large compared to the capillary length $\ell_{\gamma}\approx(d-1)\gamma\nu\Omega/(\phi^{(0)}_{\mathrm{in}}k_{\mathrm{B}}T)$. This illustrates that smaller droplets require larger concentrations of droplet material in their vicinity.
@@ 2.2.4 Dynamics of many droplets
The dynamical equations derived above for individual droplets also allow us to discuss the interaction of many droplets as long as they are well-separated. We thus consider a collection of $N_{\mathrm{d}}$} droplets described by their positions $\boldsymbol{x}_{n}$ and radii $R_{n}$ for $n=1,\ldots,{N_{\mathrm{d}}}$. These droplets are embedded in a common dilute phase, whose composition is described by a field $\phi(\boldsymbol{x})$. In the typical situation where droplets occupy only a small fraction of the entire system, it is convenient to assume that $\phi(\boldsymbol{x})$ is defined across the entire system, and that droplets provide localised perturbations [149] to this field. In this case, we describe the dynamics of the dilute phase by
@@ 2.3.1 Global parameters
Droplet size is directly influenced by the amount of droplet material in the system, $\bar{\phi}V_{\mathrm{sys}}$; see equation (2.42). Cells can thus influence whether droplets form and how large they get by producing and degrading the relevant droplet material. Droplet size also depends on the coexisting volume fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$, which in turn depend on the interaction $\chi$ and the molecular size $\Omega$; see equation (2.21). These aspects can be influenced by modifying the involved molecules, either by post-translational modifications [36, 54] or by multimerisation [172], which can induce phase separation in vitro [173] and in vivo [14, 174]. The interactions between molecules are also sensitive to environmental influences [72], including temperature [175, 176, 177], pH [178], osmotic conditions [179], crowding conditions [69], CO2 concentration [180], and electric fields [181]. Cells can exploit the phase transition associated with changes in these environmental parameters to deploy appropriate responses e.g., to control flowering in plants [182]. There is also evidence that evolution adapts genetic sequences to different temperature niches to control this behaviour [183].
@@ 3 Beyond simple liquid droplets
Cells contain an endless variety of biomolecules. For example, there are an estimated $10^{5}$ different proteins [199] present in the cell. It is thus not surprising that condensates typically consist of many different types of molecule [200, 201, 202, 55], many of which have complex individual properties themselves. Molecules can also form partially ordered structures within droplets, such as filaments [203] and porous networks [204, 205]. The complex compositional features resulting from this mixture of biomolecules affects both the bulk and interfacial properties of droplets, and also enables the coexistence of multiple different types of droplet within a cell.

### v4

@@ Abstract
  • Version of draft: November 7, 2025
@@ 2.1.1 Thermodynamics of liquids
where the first term accounts for the mixing entropy with $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$; see appendix A. We also use the molecular volume $\nu_{0}$ of the solvent to define non-dimensional molecular sizes $\Omega_{i}=\nu_{i}/\nu_{0}$. The last two terms in equation (2.1) specify enthalpic interactions using non-dimensional quantities $w_{i}$ and $\chi_{ij}$, which include solvent interactions (see equation (A.11)) and whose particular values can be determined from detailed theories [85]. Here, $w_{i}$ captures the difference of internal energy between species $i$ and the solvent, but it can also account for external potentials. Such potentials can in principle depend on space, due to, for example, gravity or gradients in a regulatory species [24], but we do not discuss this aspect further in this review. In contrast, the symmetric Flory parameters $\chi_{ij}=\chi_{ji}$ capture interactions: positive values raise the free energy and thus relate to repulsion between species $i$ and $j$, while negative values imply effective attraction. Since the interactions $\chi_{ij}$ are measured relative to solvent molecules, the diagonal terms need not vanish; see appendix A. The Flory–Huggins free energy given by equation (2.1) only describes the basic behavior of phase separating mixtures, and we discuss more complex interactions in section 3.1.
@@ 2.1.2 Kinetics of heterogeneous systems
We now derive kinetic equations based on linear non-equilibrium thermodynamics [110, 101], starting from the continuity equation, $\partial_{t}\phi_{i}+\nabla\cdot\boldsymbol{J}_{i}=s_{i}$, where $\boldsymbol{J}_{i}$ denotes spatial fluxes and the source term $s_{i}(\boldsymbol{r})$ accounts for chemical reactions as described in section 5. To separate fluxes due to hydrodynamic flows (associated with momentum transport) from diffusive fluxes driven by gradients in chemical potentials, we split $\boldsymbol{J}_{i}$ into a centre-of-mass velocity field $\boldsymbol{v}=\sum_{i=0}^{N_{\mathrm{c}}}\boldsymbol{J}_{i}$ and relative fluxes $\boldsymbol{j}_{i}=\phi_{i}(\boldsymbol{v}_{i}-\boldsymbol{v})$, where $\boldsymbol{v}_{i}=\boldsymbol{J}_{i}/\phi_{i}$ is the velocity of component $i$ in the lab frame. Since many biomolecules have mass densities close to water, we here assume equal mass density for all components, and refer the reader to ref. [101] for the general case. Hence,
@@
$\mathcal{J}_{ij}$ $=-\nu_{0}\Omega_{i}^{2}\boldsymbol{q}^{2}\sum_{l=1}^{N_{\mathrm{c}}}\Lambda_{il}\left[\frac{\partial^{2}f}{\partial\phi_{l}\partial\phi_{j}}+\kappa_{lj}\boldsymbol{q}^{2}\right]+\frac{\partial s_{i}}{\partial\phi_{j}}$ (2.15)
has only negative eigenvalues for all wave vectors $\boldsymbol{q}$. In contrast, the homogeneous state is unstable if $\mathcal{J}_{ij}$ has positive eigenvalues and the largest eigenvalue will typically determine the dominant length scale $2\pi/|\boldsymbol{q}|$ from the associated wavevector $\boldsymbol{q}$. However, since equation (2.14) is nonlinear, such a stability analysis only provides information about the initial dynamics.
@@ 2.1.3 Behaviour of binary liquids
Figure 2.1: Phase behaviour of a binary liquid. (A) Free energy density $f$ given by equation (2.16) as a function of the fraction $\phi$ (blue line) for $\Omega=1$, $\chi=2.5$, and $w=0.2$. The common-tangent construction (black line) solves equations (2.30) and thus determines the equilibrium fractions $\phi_{\mathrm{out}}$ and $\phi_{\mathrm{in}}$ (black disks) without Laplace pressure $P_{\gamma}$. Finite $P_{\gamma}$ requires two parallel tangents offset by $P_{\gamma}$ (gray lines). The inflection points (orange disks) enclose the spinodal region where $f^{\prime\prime}(\phi)<0$. (B) Exchange chemical potential $\bar{\mu}$ given by equation (2.17) as a function of $\phi$ corresponding to panel A. The blue regions of equal area indicating a Maxwell construction, equivalent to the common-tangent construction. (C) Osmotic pressure $\Pi=\nu^{-1}\phi\bar{\mu}-f$ as a function of $\phi$ corresponding to panel A. (D) $\Pi$ as a function of $\bar{\mu}$ corresponding to panel A. (E, F) Phase diagrams corresponding to panel A highlighting the equilibrium volume fractions (black binodal line) and the spinodal region (within the orange spinodal line) for a symmetric mixture ($\Omega=1$, panel E) and for large solutes ($\Omega=10$, panel F). The critical point is marked by a star and grey areas denotes the region where bi-continuous structures have lower energy than droplets in two dimensions (obtained by comparing interfacial energies of disks and stripes for equal length scales). (G, H) Grand-canonical phase diagram corresponding to panels E and F. The binodals (black lines) mark the first-order transition between dilute and dense phases. Both phases are linearly stable in the spinodal region (between orange dotted lines).
In many cases, we do not need to describe all individual components of a mixture and can instead focus on droplet material separating from a solvent. In this case, the mixture is treated as a binary fluid, and only the volume fraction $\phi$ of one component needs to be specified (${N_{\mathrm{c}}}=1$}). The local free energy density then reduces to
@@
The resulting spinodal line $\chi_{\mathrm{spin}}(\phi)$ encloses a region in phase space (figure 2.1). Within this region the homogeneous state is unstable and develops patterns at a dominant length scale $l_{\mathrm{max}}=\pi[-8\kappa/f^{\prime\prime}(\phi)]^{1/2}$ that eventually lead to phase separation, which is known as spinodal decomposition [24]. The minimum of the curve $\chi_{\mathrm{spin}}(\phi)$ corresponds to the critical point,
@@ 2.1.4 Interfaces and phase morphologies
In the absence of chemical reactions ($s=0$), the dynamics described by equation (2.23) result in equilibrium states that minimise the free energy $F$ given by equation (2.22). Figure 2.2A shows that the corresponding equilibrium profile $\phi_{\mathrm{eq}}(x)$ smoothly interpolates between the fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ of the two co-existing phases. Denoting by $x$ a coordinate perpendicular to the interface positioned at $x=0$, we approximate the profile as
$\phi_{\mathrm{eq}}(x)$ $\approx\frac{\phi^{(0)}_{\mathrm{in}}+\phi^{(0)}_{\mathrm{out}}}{2}+\frac{\phi^{(0)}_{\mathrm{in}}-\phi^{(0)}_{\mathrm{out}}}{2}\tanh\left(\frac{2x}{0pt}\right)$ (2.24)
and determine the interfacial width $0pt$ below. This expression is exact for simple polynomial free energies [134, 24], and it captures the sigmoidal profile of more general cases. For the free energy density given by equation (2.16) with symmetric molecular sizes ($\Omega=1$), the interfacial width $0pt$ is approximately
@@
which is valid for weak segregation ($\chi\gtrsim 2$); see appendix C. This expression indicates that stronger segregation, i.e., larger $\phi^{(0)}_{\mathrm{in}}-\phi^{(0)}_{\mathrm{out}}$, implies narrower interfaces.
In the typical case where the region of the interface is thin compared to the linear size of the adjacent phases, one can use a thin-interface approximation and replace the sigmoidal profile by a step function. The excess free energy, which is not accounted for by the bulk phases, defines the surface energy [104]
@@
where $\bar{\mu}_{\mathrm{eq}}$ is the exchange chemical potential and $\Pi_{\mathrm{eq}}=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{out}}-f(\phi^{(0)}_{\mathrm{out}})=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{in}}-f(\phi^{(0)}_{\mathrm{in}})$ denotes the associated osmotic pressure; see appendix C. The second line assumes a monotonic profile $\phi_{\mathrm{eq}}(x)$, which may not be valid in complex mixtures [135]. However, $\phi_{\mathrm{eq}}(x)$ is always monotonic in binary mixtures, and in the symmetric case $\Omega=1$, we find that (appendix C)
$\gamma$ $\approx\sqrt{\frac{k_{\mathrm{B}}T\kappa}{2\nu}}\cdot\frac{\phi^{(0)}_{\mathrm{out}}-\phi^{(0)}_{\mathrm{in}}}{3}\ln\left(4\phi^{(0)}_{\mathrm{in}}\phi^{(0)}_{\mathrm{out}}\right)\;$, (2.27)
@@
The thin-interface approximation simplifies the analysis of equilibrium states in binary liquids: material conservation fixes the total volume of the dilute and dense phases, so their shape is the only degree of freedom. The total free energy thus comprises a constant term covering the bulk contributions and an interfacial term, $\gamma A$, where $A$ is the total interfacial area. Equilibrium states thus minimise $A$, which implies constant mean curvature of interfaces [138]. Consequently, surface energies can be interpreted as a mechanical surface tension [139], explaining why the minority phase favours spherical shapes. Conversely, if both phases occupy similar volumes, bicontinuous structures, such as cylinders and flat lamella, can emerge; see figure 2.2B. These structures predominate in a region of phase space that overlaps significantly with the spinodal region; see figure 2.1E,F. However, these two phenomena are distinct, so droplets can for instance emerge from spinodal decomposition. Bicontinuous structures are rarely observed in biological cells, and so we exclusively focus on droplets in this review.
@@ 2.2.1 Local equilibrium at the interface
where $c_{\mathrm{in}}=\phi^{(0)}_{\mathrm{in}}/\nu\Omega$ is the number concentration inside the droplet and the dilute phase is approximated as an ideal solution. When only surface tension acts on the droplet, $\Delta P=P_{\gamma}$, we find the Gibbs-Thomson relation [142]
@@
assuming the radius $R$ is large compared to the capillary length $\ell_{\gamma}\approx(d-1)\gamma\nu\Omega/(\phi^{(0)}_{\mathrm{in}}k_{\mathrm{B}}T)$. This illustrates that smaller droplets require larger concentrations of droplet material in their vicinity to be in equilibrium with their surroundings.
@@ 2.2.4 Dynamics of many droplets
The dynamical equations derived above for individual droplets also allow us to discuss the interaction of many droplets as long as they are well-separated. We thus consider a collection of $N_{\mathrm{d}}$} droplets described by their positions $\boldsymbol{x}_{n}$ and radii $R_{n}$ for $n=1,\ldots,{N_{\mathrm{d}}}$. These droplets are embedded in a common dilute phase, whose composition is described by a field $\phi(\boldsymbol{x})$. In the typical situation where droplets occupy only a small fraction of the entire system, it is convenient to assume that $\phi(\boldsymbol{x})$ is defined across the entire system, and that droplets provide localised perturbations to this field [149]. In this case, we describe the dynamics of the dilute phase by
@@ 2.3.1 Global parameters
Droplet size is directly influenced by the amount of droplet material in the system, $\bar{\phi}V_{\mathrm{sys}}$; see equation (2.42). Cells can thus influence whether droplets form and how large they get by producing and degrading the relevant droplet material. Droplet size also depends on the coexisting volume fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$, which in turn depend on the interaction $\chi$ and the molecular size $\Omega$; see equation (2.21). These aspects can be influenced by modifying the involved molecules, either by post-translational modifications [36, 54] or by multimerisation [172], which can induce phase separation in vitro [173] and in vivo [14, 174]. The interactions between molecules are also sensitive to environmental influences [72], including temperature [175, 176, 177], pH [178], osmotic conditions [179], crowding conditions [69], CO2 concentration [180], and electric fields [181]. Cells can exploit the phase transition associated with changes in these environmental parameters to deploy appropriate responses, e.g., to control flowering in plants [182]. There is also evidence that evolution adapts genetic sequences to different temperature niches to control this behaviour [183].
@@ 3 Beyond simple liquid droplets
Cells contain an endless variety of biomolecules. For example, there are an estimated $10^{5}$ different proteins [199] present in the cell. It is thus not surprising that condensates typically consist of many different types of molecules [200, 201, 202, 55], many of which have complex individual properties themselves. Molecules can also form partially ordered structures within droplets, such as filaments [203] and porous networks [204, 205]. The complex compositional features resulting from this mixture of biomolecules affects both the bulk and interfacial properties of droplets, and also enables the coexistence of multiple different types of droplet within a cell.